# Joint Distillation for Fast Likelihood Evaluation and Sampling in Flow-based Models

### v1

@@ 4 Related works
Likelihood computation in diffusion and flow models. While diffusion and flow-based models excel at sample generation, their likelihood evaluation remains computationally expensive. Discrete-time diffusion models compute likelihoods through variational bounds requiring hundreds of NFEs (ho2020denoising; nichol2021improved). Continuous formulations enable exact likelihood via the probability flow ODE (song2020score) but require numerical integration with 100-1000 NFEs. Prior research has explored various techniques to improve likelihood estimation (grathwohl2018ffjord; song2021maximum), but they still require many NFEs for accurate evaluation. In parallel, normalizing flows (rezende2015variational) offer exact and tractable log-likelihoods by using specialized network architecture designs and change-of-variables formula. Many recent efforts (zhai2024normalizing; ho2019flow++; chen2019residual) aim at scaling up the normalizing flow principles through better dequantization and and advance architectures.
Accelerating sampling in flow-based models. Reducing sampling costs has been a major focus in diffusion and flow matching research. Advanced ODE solvers (karras2022elucidating; lu2022dpm) leverage the semi-linear structure of the probability flow ODE to reduce discretization error. Distillation methods (salimans2022progressive; sauer2024adversarial) and self-distillation models (song2023improved; zhou2025inductive) provide alternative solutions by training student models to match teacher trajectories or training to perform self-bootstrapping with fewer steps. In particular, consistency models (song2023consistency) and consistency trajectory models (kim2023consistency) learn direct mappings from any point along the trajectory to data.
Flow map-based methods. Flow maps provide a general framework to model the solution operator of ODEs, enabling direct prediction of integrated trajectories (kim2023consistency; boffi2025buildconsistencymodellearning; boffi2025flowmapmatchingstochastic). Recent works exploit this structure for few-step sampling. For example, as we have shown above, shortcut models (frans2024one) imposes semigroup property to learn the flow maps while MeanFlow (geng2025mean) and Align Your Flow (sabour2025align) enforce Eulerian conditions. While these methods successfully reduce sampling to less than 10 NFEs, they either abandon likelihood computation entirely or still require full trajectory integration for likelihood evaluation.
@@ 5.1 Setups
We empirically verify the effectiveness of our method on image datasets CIFAR-10 (cifar10) and ImageNet $64\times 64$ (imagenet). We evaluate the sample quality using Fréchet Inception Distance (FID) (heusel2017gans) on 50K generated images. The negative log-likelihood (NLL) is measured in bits per dimension (BPD) on the entire test set of CIFAR-10 and a randomly sampled 10K subset of the ImageNet test set. We compare our method against flow matching (lipman2022flow), shortcut models (frans2024one) and MeanFlow (geng2025mean) as baselines, and augment the later two for joint distillation. All models are unconditionally trained on both datasets. We use $1,2,4,8$ Euler steps for sampling and likelihood evaluation in both experiments. Implementation details about our method and the baselines can be found in Appendix C.
@@ Appendix A Characterization of the joint flow map
We first recall the simple tangent identity denoted the tangent condition by boffi2025buildconsistencymodellearning, also leveraged by (kim2023consistency; frans2024one; geng2025mean), which allows us to recover the instantaneous velocity and divergence in the $s\to t$ limit:
@@
The proof follows by application of Lemma 2.1 from boffi2025buildconsistencymodellearning. ∎
@@
The proof follows by application of Proposition 2.2 from boffi2025buildconsistencymodellearning applied to the joint system. ∎
Proposition˜A.2 gives three characterizations of the joint flow map that may be used to devise few-step flow-based training algorithms. In each case, by definition of Eq.˜A.3, the $X$ block reduces to the flow map characterizations introduced in boffi2025flowmapmatchingstochastic; boffi2025buildconsistencymodellearning for the sampling system. The second block for the likelihood dynamics is new, which we focus on now to instantiate the resulting equations.
@@
Setting $r=\frac{1}{2}(t+s)$ recovers a continuous limit of shortcut models, as shown in boffi2025buildconsistencymodellearning. In this case, Eq.˜A.18 becomes
@@ Appendix B Additional Theoretical Analysis
MeanFlow (geng2025mean) directly enforces the Eulerian condition with MeanFLow identity.
@@ Appendix C Implementation Details
Shortcut Model. We reimplement the shortcut model ( frans2024one) for CIFAR-10. Following their method, we use discrete timesteps and set $1/128$ as the smallest unit of time for approximating the ODE. We consider 8 possible shortcut lengths ranging from $(1,1/2,\ldots,1/128)$. Similarly, we divide the batch into $3/4$ for flow-matching training and $1/4$ for self-consistency training. However, instead of DiT, we use a U-Net backbone with configurations in Table 3. We train for 100k iterations on CIFAR-10 and sample with 1, 2, 4, and 8 standard Euler steps. Additionally, we follow the original paper and parametrize the model to take $x_{t}$, $t$ and sts-t as inputs.
@@
BPD. We use the same method as lipman2022flow to compute the BPD. To compute BPD with F2D2 models, we discretize the interval from 1 to 0 into equal segments, while the uniform step size serves as the second time input.
2D Checkerboard. We follow the experimental setup of boffi2025buildconsistencymodellearning. The model is a 4-layer MLP with 512 hidden units per layer and GELU activations. Both the flow-matching baseline and the Shortcut model are trained for 150k iterations with a batch size of 100k and an initial learning rate of $10^{-3}$, using a square-root decay schedule after 35k steps. For the Shortcut model, each batch is divided in a $3:1$ ratio between the flow-matching loss and the self-distillation loss. For Shortcut-F2D2, we extend the output layer with an additional dimension and initialize training from the pretrained Shortcut model, running an extra 27k iterations.
@@ Appendix D Additional Ablations
The experiment in Table 6 opens the door to many intriguing future research directions that can potentially further improve the performance. These include more fine-grained loss weighting scheme and developing training schedules similar to the ones proposed in frans2024one and geng2025mean, where the flow matching objective and the self-consistency objective are optimized in separate portions of the training.

### v2

@@ 4 Related works
Likelihood computation in diffusion and flow models. While diffusion and flow-based models excel at sample generation, their likelihood evaluation remains computationally expensive. Discrete-time diffusion models compute likelihoods through variational bounds requiring hundreds of NFEs (Ho et al., 2020; Nichol & Dhariwal, 2021). Continuous formulations enable exact likelihood via the probability flow ODE (Song et al., 2020) but require numerical integration with 100-1000 NFEs. Prior research has explored various techniques to improve likelihood estimation (Grathwohl et al., 2018; Song et al., 2021), but they still require many NFEs for accurate evaluation. In parallel, normalizing flows (Rezende & Mohamed, 2015) offer exact and tractable log-likelihoods by using specialized network architecture designs and change-of-variables formula. Many recent efforts (Zhai et al., 2024; Ho et al., 2019; Chen et al., 2019) aim at scaling up the normalizing flow principles through better dequantization and and advance architectures.
Accelerating sampling in flow-based models. Reducing sampling costs has been a major focus in diffusion and flow matching research. Advanced ODE solvers (Karras et al., 2022; Lu et al., 2022) leverage the semi-linear structure of the probability flow ODE to reduce discretization error. Distillation methods (Salimans & Ho, 2022; Sauer et al., 2024) and self-distillation models (Song & Dhariwal, 2023; Zhou et al., 2025) provide alternative solutions by training student models to match teacher trajectories or training to perform self-bootstrapping with fewer steps. In particular, consistency models (Song et al., 2023) and consistency trajectory models (Kim et al., 2023) learn direct mappings from any point along the trajectory to data.
Flow map-based methods. Flow maps provide a general framework to model the solution operator of ODEs, enabling direct prediction of integrated trajectories (Kim et al., 2023; Boffi et al., 2025a; b). Recent works exploit this structure for few-step sampling. For example, as we have shown above, shortcut models (Frans et al., 2024) imposes semigroup property to learn the flow maps while MeanFlow (Geng et al., 2025) and Align Your Flow (Sabour et al., 2025) enforce Eulerian conditions. While these methods successfully reduce sampling to less than 10 NFEs, they either abandon likelihood computation entirely or still require full trajectory integration for likelihood evaluation.
@@ 5.1 Setups
We empirically verify the effectiveness of our method on image datasets CIFAR-10 (Krizhevsky et al., ) and ImageNet $64\times 64$ (Deng et al., 2009). We evaluate the sample quality using Fréchet Inception Distance (FID) (Heusel et al., 2017) on 50K generated images. The negative log-likelihood (NLL) is measured in bits per dimension (BPD) on the entire test set of CIFAR-10 and a randomly sampled 10K subset of the ImageNet test set. We compare our method against flow matching (Lipman et al., 2022), shortcut models (Frans et al., 2024) and MeanFlow (Geng et al., 2025) as baselines, and augment the later two for joint distillation. All models are unconditionally trained on both datasets. We use $1,2,4,8$ Euler steps for sampling and likelihood evaluation in both experiments. Implementation details about our method and the baselines can be found in Appendix C.
@@ Appendix A Characterization of the joint flow map
We first recall the simple tangent identity denoted the tangent condition by Boffi et al. (2025a), also leveraged by (Kim et al., 2023; Frans et al., 2024; Geng et al., 2025), which allows us to recover the instantaneous velocity and divergence in the $s\to t$ limit:
@@
The proof follows by application of Lemma 2.1 from Boffi et al. (2025a). ∎
@@
The proof follows by application of Proposition 2.2 from Boffi et al. (2025a) applied to the joint system. ∎
Proposition˜A.2 gives three characterizations of the joint flow map that may be used to devise few-step flow-based training algorithms. In each case, by definition of Eq.˜A.3, the $X$ block reduces to the flow map characterizations introduced in Boffi et al. (2025b; a) for the sampling system. The second block for the likelihood dynamics is new, which we focus on now to instantiate the resulting equations.
@@
Setting $r=\frac{1}{2}(t+s)$ recovers a continuous limit of shortcut models, as shown in Boffi et al. (2025a). In this case, Eq.˜A.18 becomes
@@ Appendix B Additional Theoretical Analysis
MeanFlow (Geng et al., 2025) directly enforces the Eulerian condition with MeanFLow identity.
@@ Appendix C Implementation Details
Shortcut Model. We reimplement the shortcut model ( Frans et al. (2024)) for CIFAR-10. Following their method, we use discrete timesteps and set $1/128$ as the smallest unit of time for approximating the ODE. We consider 8 possible shortcut lengths ranging from $(1,1/2,\ldots,1/128)$. Similarly, we divide the batch into $3/4$ for flow-matching training and $1/4$ for self-consistency training. However, instead of DiT, we use a U-Net backbone with configurations in Table 3. We train for 100k iterations on CIFAR-10 and sample with 1, 2, 4, and 8 standard Euler steps. Additionally, we follow the original paper and parametrize the model to take $x_{t}$, $t$ and sts-t as inputs.
@@
BPD. We use the same method as Lipman et al. (2022) to compute the BPD. To compute BPD with F2D2 models, we discretize the interval from 1 to 0 into equal segments, while the uniform step size serves as the second time input.
2D Checkerboard. We follow the experimental setup of Boffi et al. (2025a). The model is a 4-layer MLP with 512 hidden units per layer and GELU activations. Both the flow-matching baseline and the Shortcut model are trained for 150k iterations with a batch size of 100k and an initial learning rate of $10^{-3}$, using a square-root decay schedule after 35k steps. For the Shortcut model, each batch is divided in a $3:1$ ratio between the flow-matching loss and the self-distillation loss. For Shortcut-F2D2, we extend the output layer with an additional dimension and initialize training from the pretrained Shortcut model, running an extra 27k iterations.
@@ Appendix D Additional Ablations
The experiment in Table 6 opens the door to many intriguing future research directions that can potentially further improve the performance. These include more fine-grained loss weighting scheme and developing training schedules similar to the ones proposed in Frans et al. (2024) and Geng et al. (2025), where the flow matching objective and the self-consistency objective are optimized in separate portions of the training.